OpenStudy (anonymous):

how the force between two bodies is affected if the distance between them is tripled? plz.....answer fast urgent....pls.

OpenStudy (anonymous):

inverse sqaure relation

OpenStudy (amistre64):

nothing should be THAT urgent ... What is the formula for gravitational force between to objects?

OpenStudy (anonymous):

\[f=\frac{ GMm }{ d ^{2} }\]

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

now take the distance and triple it; then compare the 2 results \[f=\frac{ GMm }{ (3d) ^{2} }\]

OpenStudy (anonymous):

First thing - @basith think: does the force decay (weaken) or increase when the distance is increasing ?...

OpenStudy (anonymous):

Aand - this process is QUADRATIC. So together with first fact it yields what conclusion ?

OpenStudy (amistre64):

\[f_1=\frac{ GMm }{ d ^{2} }\to\ f_3=\frac{1}{9}*\frac{ GMm }{ d ^{2} }\]

OpenStudy (anonymous):

\[thats,=\frac{ 1 }{ 9 }F\]

OpenStudy (anonymous):

then what?

OpenStudy (amistre64):

then write that down and move to the next question?

OpenStudy (amistre64):

when the distance is tripled, the force is one-ninth of what it was ...

OpenStudy (anonymous):

k thnx